W lo Pharmacology B) Has she ever received gentamicin? C) Has she ever received ampicillin D) Has she ever received cephal
Answer #1

ANSWER.

44) D) Nurse need to be alert for any medication that cause allergic reaction

     to reduce the risk of adverse effect during drug therapy.

45) Calculation of drops per minute.

          Formula

          Drops per minute(gtt/min) = Volume(ml) /Time(min) * Drop factor(gtt/ml).

      Here

           Volume(ml) = 500 ml.

           Time(min) = 480 min (8 hr).

           Drop factor = 15 drops per ml

       As per above formula

            Drops per minute(gtt/min) = 500 / 480 * 15

                                                        = 1.04 * 15

                                                        = 15.6 gtt/min.

46) Calculation of volume(ml) of drug required.

             Formula

                Volume(ml) = Desired dose/Dose in hand * Quantity.

         Here

              Desired dose(D) = 7500 units.

              Dose in hand (H) = 10000 units.

                   Quantity (Q) = 1 ml.

    As per above formula

                Volume = D/H * Q

                             = 7500/10000 * 1

                             = 0.75 * 1

                          B = 0.75 ml.

47) C) Nursing diagnosis for an older client live independently has just prescribed a

     drug that cause sedation is Risk for injury.